# From Tokens to Thoughts: How LLMs and Humans Trade Compression for Meaning

### v3

@@ A.3 Additional Clustering Metrics
Figure 3: LLM-derived Clusters Show Above-Chance Alignment with Human Conceptual Categories. Normalized Mutual Information (NMI) between human-defined categories and clusters from LLM embeddings. Results are averaged over three psychological datasets. All models perform significantly better than random clustering. BERT’s performance is notably strong.
Figure 4: LLM-derived Clusters Show Above-Chance Alignment with Human Conceptual Categories. Adjusted Rand Index (ARI) between human-defined categories and clusters from LLM embeddings. Results are averaged over three psychological datasets. All models perform significantly better than random clustering. BERT’s performance is notably strong.
@@
A.4 Detailed AMI Scores per Model and Dataset
Table LABEL:tab:ami_detailed provides a more granular view of the AMI scores for each LLM across the three individual psychological datasets.
@@ A.4 Detailed AMI Scores per Model and Dataset
A.5 Correlation between Human Typicality Judgments and LLM Internal Cluster Geometry
Model Dataset Correlation (Spearman $\rho$)
@@ A.5 Correlation between Human Typicality Judgments and LLM Internal Cluster Geometry
Qwen1.5-72B -0.237 -0.049 -0.016
Llama-3-70B -0.124∗∗ -0.085 0.016
Llama-3.1-70B -0.125∗∗ -0.084 0.015
Qwen1.5-32B -0.051 -0.064∗∗ 0.007
gemma-2-27b -0.166 -0.116 0.038
Qwen1.5-14B -0.197 -0.052 -0.029
phi-4 -0.061 -0.044 0.025
gemma-2-9b -0.282 -0.074 0.117
Llama-3.1-8B -0.184 -0.075 -0.058
Llama-3-8B -0.162 -0.073 -0.053
Mistral-7B-v0.3 0.015 -0.112 0.040
Qwen2-7B -0.021 -0.105 -0.008
Qwen2.5-7B 0.033 -0.066 -0.030
gemma-7b -0.135 -0.047∗∗ 0.010
Llama-3.2-3B -0.007 0.000 0.001
phi-2 0.049 -0.108∗∗ -0.001
gemma-2b -0.176 -0.055 0.052
gemma-2-2b -0.283 -0.107 0.117
Qwen1.5-1.8B -0.106 -0.085 0.021
Qwen2.5-1.5B -0.003 -0.035 0.015
phi-1.5 0.134 -0.134 0.007
phi-1 -0.219 -0.138 0.013
Llama-3.2-1B -0.062 -0.004∗∗ -0.003
Qwen1.5-0.5B -0.122 -0.004 -0.001
Qwen2-0.5B -0.044 0.009 -0.009
Qwen2.5-0.5B -0.018 -0.009 -0.007
roberta-large 0.088 -0.047 -0.074
bert-large-uncased -0.427 -0.198∗∗ 0.206∗∗
deberta-large 0.016 -0.042 -0.023
Table 2: Correlation between Human Typicality Judgments and LLM Internal Cluster Geometry. Spearman rank correlations between human-rated psychological typicality/distance (higher human scores = less typical/more distant) and item-to-centroid cosine similarity (higher similarity = more central to LLM cluster). Negative correlations suggest alignment. ${}^{**}p<0.05$.
A.6 Typicality and Cosine Similarity [RQ2]
@@
Figure 6: Weak and Mostly Non-Significant Spearman Correlation Values Between Human Typicality Judgments and LLM Cosine Similarity Indicating Different Structure Representing Concepts. Mean Spearman correlation values across the models belonging to the same family and across the three datasets.
A.7 Theoretical Extreme Case Exploration for $\mathcal{L}$
(Content from your original Appendix Section A: “Theoretical Extreme Case Exploration“ would go here, ensuring it refers to $\mathcal{L}$ as defined in Equation 4).
@@ A.7 Theoretical Extreme Case Exploration for $\mathcal{L}$
A.8 Compression Figures
Figure 7 shows the mean cluster entropy ($S_{\alpha}$) versus the number of clusters ($K$) aggregated across the different LLM families, compared against human-defined categories (represented as distinct points or lines at their fixed $K$ values from the datasets). Higher entropy values indicate less compressed or more diverse clustering.
@@ A.8 Compression Figures
Figure 7: Human Conceptual Categories Exhibit Higher Mean Entropy than LLM-Derived Clusters. Mean cluster entropy ($S_{\alpha}$) versus the number of clusters ($K$) for various LLMs, compared against human-defined categories (represented as distinct points or lines at their fixed $K$ values from the datasets). Higher entropy values indicate less compressed or more diverse clusterings.
Figure 8: LLMs Achieve a More “Optimal” Compression-Meaning Trade-off by the $\mathcal{L}$ Measure. IB-RDT objective ($\mathcal{L}$) vs. $K$. Lower $\mathcal{L}$ indicates a more optimal balance between compression ($I(X;C)$) and semantic fidelity (distortion). Human categories (fixed $K$) show higher $\mathcal{L}$ values.

### v4

@@ A.8 Additional Clustering Metrics
Figure 12: LLM-derived Clusters Show Above-Chance Alignment with Human Conceptual Categories. Normalized Mutual Information (NMI) between human-defined categories and clusters from static LLM embeddings. Results are averaged over three psychological datasets. All models perform significantly better than random clustering. BERT’s performance is notably strong.
Figure 13: LLM-derived Clusters Show Above-Chance Alignment with Human Conceptual Categories. Adjusted Rand Index (ARI) between human-defined categories and clusters from LLM static embeddings. Results are averaged over three psychological datasets. All models perform significantly better than random clustering. BERT’s performance is notably strong.
@@
A.9 Detailed AMI Scores per Model and Dataset
Table LABEL:tab:ami_detailed provides a more granular view of the static AMI scores for each LLM across the three individual psychological datasets.
@@ A.9 Detailed AMI Scores per Model and Dataset
A.10 Correlation between Human Typicality Judgments and LLM Internal Cluster Geometry
The following tables present the Spearman correlation coefficients ($\rho$) between human typicality judgments and LLM internal representations across different analysis approaches:
Table 3: Static analysis correlations using embeddings from the E matrix. This approach captures the baseline semantic relationships between items and categories without contextual processing.
Table 4: Peak AMI layer analysis correlations using contextual embeddings from the layer that maximized AMI scores (as identified in RQ1). This approach leverages the optimal layer for semantic clustering to assess fine-grained semantic fidelity.
Both tables present correlations across three cognitive science datasets: Rosch (1973), Rosch (1975), and McCloskey (1978), with asterisks (*) indicating statistically significant correlations ($p<0.05$). The modest correlation values across most models suggest limited alignment between LLM internal representations and human-perceived semantic nuances.
Model Dataset Correlation (Spearman $\rho$) - Static Layer
@@ A.10 Correlation between Human Typicality Judgments and LLM Internal Cluster Geometry
Deberta large (304M) 0.144 0.107* 0.075
Bert large (340M) 0.378* 0.275* 0.250*
Roberta large (355M) 0.005 0.038 -0.029
Gemma (2B) 0.069 0.078 0.007
Gemma 2 (2B) 0.236 0.119* 0.147*
Gemma (7B) 0.131 0.100* 0.007
Gemma 2 (9B) 0.280 0.135* 0.199*
Gemma 2 (27B) 0.112 0.122* 0.161*
Qwen1.5 (0.5B) 0.175 0.076 0.096*
Qwen2 (0.5B) 0.238 0.041 0.040
Qwen2.5 (0.5B) 0.212 0.027 0.037
Qwen2.5 (1.5B) 0.141 0.086* 0.078
Qwen1.5 (1.8B) 0.172 0.134* 0.154*
Qwen2 (7B) 0.036 0.087* 0.040
Qwen1.5 (14B) 0.154 0.086* 0.108*
Qwen1.5 (32B) -0.032 0.100* 0.081
Qwen2.5 (32B) 0.035 0.105* 0.084
Mistral v0.3 (7B) 0.076 0.152* -0.009
Llama 3.2 (1B) 0.301* 0.056 0.039
Llama 3 (8B) -0.002 0.099* 0.080
Llama 3.1 (8B) 0.004 0.108* 0.081
Llama 3 (70B) 0.148 0.161* 0.155*
Llama 3.1 (70B) 0.122 0.161* 0.155*
Phi 1 (1.42B) 0.071 0.052 0.054
Phi 1.5 (1.42B) -0.088 0.079 0.018
Phi 2 (2.78B) -0.056 0.044 0.024
Phi 4 (14.7B) 0.079 0.086* 0.097*
T5 Large (770M) 0.235 0.259* 0.178*
GPT-2 Medium (355M) -0.032 0.063 -0.017
ViT-B/32 Text (63.1M) 0.527* 0.315* 0.286*
ViT-B/16 Text (63.1M) 0.528* 0.289* 0.278*
Word2Vec (300D) 0.442* 0.349* 0.437*
Glove (300D) 0.315* 0.333* 0.350*
Table 3: Correlation between Human Typicality Judgments and LLM Internal Cluster Geometry. Spearman static-layer rank correlations between human-rated psychological typicality/distance (higher human scores = less typical/more distant) and item-to-centroid cosine similarity (higher similarity = more central to LLM cluster). $^{*}p<0.05$.
Model Dataset Correlation (Spearman $\rho$) - Peak AMI Layer
Rosch (1973) Rosch (1975) McCloskey (1978)
Deberta large (304M) 0.277 0.107* 0.126*
Bert large (340M) -0.120 0.148* 0.026
Roberta large (355M) -0.011 0.038 -0.022
Gemma (2B) 0.127 0.092* 0.039
Gemma 2 (2B) 0.034 0.139* 0.103*
Gemma (7B) 0.004 -0.050 0.088
Gemma 2 (9B) 0.047 0.110* 0.098*
Gemma 2 (27B) -0.135 0.090* -0.101*
Qwen1.5 (0.5B) -0.025 -0.064 0.122*
Qwen2 (0.5B) 0.090 0.064 -0.077
Qwen2.5 (0.5B) -0.012 0.121* 0.072
Qwen2.5 (1.5B) -0.114 0.084* -0.028
Qwen1.5 (1.8B) 0.092 -0.044 -0.004
Qwen2 (7B) 0.004 0.049 -0.087
Qwen1.5 (14B) 0.032 0.091* 0.043
Qwen1.5 (32B) 0.026 0.082 0.111*
Qwen2.5 (32B) 0.045 0.079 -0.060
Mistral v0.3 (7B) 0.013 0.039 0.107*
Llama 3.2 (1B) 0.063 0.094* 0.070
Llama 3 (8B) -0.045 0.138* 0.084
Llama 3.1 (8B) -0.096 0.130* 0.087
Llama 3 (70B) -0.108 0.023 0.050
Phi 1 (1.42B) 0.216 0.185* -0.029
Phi 1.5 (1.42B) 0.162 0.098* 0.015
Phi 2 (2.78B) 0.325* 0.142* -0.033
Phi 4 (14.7B) 0.079 0.129* 0.007
T5 Large (770M) 0.219 0.226* 0.282*
GPT-2 Small (117M) -0.046 0.118* 0.010
GPT-2 Medium (355M) 0.077 0.109* -0.029
ViT-B/32 Text (63.1M) 0.128 0.055 0.152*
ViT-B/16 Text (63.1M) 0.089 0.086* 0.127*
Word2Vec (300D) 0.442* 0.349* 0.437*
Glove (300D) 0.315* 0.333* 0.350*
Table 4: Correlation between Human Typicality Judgments and LLM Internal Cluster Geometry. Spearman peak-AMI layer rank correlations between human-rated psychological typicality/distance (higher human scores = less typical/more distant) and item-to-centroid cosine similarity (higher similarity = more central to LLM cluster). $^{*}p<0.05$.
A.11 Typicality and Cosine Similarity [RQ2]
@@
Figure 15: Weak and Mostly Non-Significant Spearman Correlation Values Between Human Typicality Judgments and LLM Cosine Similarity Indicating Different Structure Representing Concepts. Mean Static Layer Spearman correlation values across the models belonging to the same family and across the three datasets.
Figure 16: Weak and Mostly Non-Significant Spearman Correlation Values Between Human Typicality Judgments and LLM Cosine Similarity Indicating Different Structure Representing Concepts. Mean Peak AMI Layer Spearman correlation values across the models belonging to the same family and across the three datasets.
A.12 Theoretical Extreme Case Exploration for $\mathcal{L}$
@@
A.13 Compression Figures
@@
Figure 17: Static Embeddings Achieve a more “Optimal” Compression-Meaning Trade-off by the $\mathcal{L}$ Measure. IB-RDT objective ($\mathcal{L}$) vs. $K$ across all datasets. Lower $\mathcal{L}$ indicates a more optimal balance between compression ($I(X;C)$) and semantic fidelity (distortion). Static embeddings consistently achieve lower $\mathcal{L}$ values than both human categories and contextual embeddings. The plots correspond to the three datasets in the following order: Rosch [1973a], Rosch [1975], McCloskey and Glucksberg [1978].
Figure 18: Contextual Embeddings Achieve Better-than-Human Compression-Meaning Trade-off by the $\mathcal{L}$ Measure. IB-RDT objective ($\mathcal{L}$) vs. $K$ across all datasets. Lower $\mathcal{L}$ indicates a more optimal balance between compression ($I(X;C)$) and semantic fidelity (distortion). Contextual embeddings outperform human categories but achieve higher $\mathcal{L}$ values than static embeddings. The plots correspond to the three datasets in the following order: Rosch [1973a], Rosch [1975], McCloskey and Glucksberg [1978].
A.14 $\mathcal{L}$ objective vs. downstream task performance
Analysis of 13 instruction-tuned models across 5 families (Qwen, Llama, Gemma, Phi, Mistral; see Table 5 for results) indicates no statistical significance (r=0.202r=-0.202, $p=0.508$). This finding suggests that while the $\mathcal{L}$ objective successfully identifies models that compress semantic categories more effectively, this compression ability does not directly translate to improved performance on standard NLP benchmarks. The lack of correlation implies that concept compression and benchmark accuracy represent distinct aspects of model capability, with the former capturing semantic organization efficiency and the latter measuring general knowledge and reasoning abilities. We specifically chose instruction-tuned models to ensure fair comparison on MMLU, as base models would likely perform poorly on this instruction-following benchmark. While our analysis covers a diverse range of model families and sizes, this represents a subset of available models due to the limited availability of instruction-tuned variants.
Model Size $\mathcal{L}$ Score MMLU Score
Qwen2-0.5B-Instruct 494M 1.930 0.433
Qwen2.5-0.5B-Instruct 494M 1.982 0.469
Llama-3.2-1B-Instruct 1.2B 1.876 0.454
Qwen2.5-1.5B-Instruct 1.5B 2.071 0.597
Gemma-2B-IT 2.5B 2.382 0.366
Gemma-2-2B-IT 2.6B 2.263 0.565
Phi-4-mini-instruct 3.8B 1.905 0.678
Mistral-7B-Instruct-v0.3 7.2B 1.714 0.603
Qwen2-7B-Instruct 7.6B 2.319 0.700
Meta-Llama-3-8B-Instruct 8.0B 1.348 0.647
Llama-3.1-8B-Instruct 8.0B 1.320 0.679
Gemma-7B-IT 8.5B 2.372 0.512
Gemma-2-9B-IT 9.2B 2.467 0.723
Table 5: No correlation between $\mathcal{L}$ objective scores and MMLU scores across different sizes and families. Table displays $\mathcal{L}$ objective values vs. MMLU scores for 13 instruction-tuned models. Correlation measured (r=0.202r=-0.202, $p=0.508$).